The complement of an angle is one-fifth of the measure of the angle. Find the measure of the angle.
Answer by stanbon(75887)

The complement of an angle is one-fifth of the measure of the angle. Find the measure of the angle.
----------------------
Let the measure of the angle be "x".
It's compliment is "90-x"
--------------
EQUATION:
90-x = (1/5)x
(6/5)x = 90
6x = 450
x = 75
----------
Cheers,
Stan H.